Math 34 Differential Equations Quiz 1 Sept 1, 2010 SHOW ALL WORK

1.) Give an example of a linear differential equation:

2.) Give an example of a non-linear differential equation:

3.) From section 1.4, name one mathematician who studied differential equations OR one application (but not related to gravity) of differential equations.

4.) Circle T for True or F for False:

4a.) Numerical approximations for solutions to differential equations are often needed as the solutions to most differential equations cannot be expressed algebraically.

T F

4b.) If a computer is used to find a numerical approximation to a differential equation, then we know the equation has at least one solution.

T F